OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -3/8 and passing through (5, -4)

OpenStudy (cggurumanjunath):

y=mx+c--- standard form the equation of the given line (x,y)=(5,-4)

OpenStudy (cggurumanjunath):

m=-3/8

OpenStudy (cggurumanjunath):

y=mx+c becomes -4=(-3/8)(5)+c

OpenStudy (cggurumanjunath):

now can u find constant c ?

OpenStudy (cggurumanjunath):

@Power2Knowledge ?

OpenStudy (cggurumanjunath):

m-slope and c is constant !

OpenStudy (anonymous):

Im new to these equations

OpenStudy (anonymous):

hello

OpenStudy (cggurumanjunath):

y=mx+c--- standard form the equation of the given line is this clear ? are u aware of this equation !

OpenStudy (anonymous):

a little.

OpenStudy (cggurumanjunath):

y=mx+c--- standard form the equation of the given line

OpenStudy (anonymous):

so my answer is -4=(-3/8)(5)+c?

OpenStudy (cggurumanjunath):

now can u find constant c !

OpenStudy (anonymous):

how do i find that?

OpenStudy (cggurumanjunath):

-4=(-3/8)(5)+c

OpenStudy (cggurumanjunath):

-4=(-15/8)+C

OpenStudy (cggurumanjunath):

now can u find constant C !

OpenStudy (anonymous):

40?

OpenStudy (cggurumanjunath):

HOW ?

OpenStudy (anonymous):

i multiplied 5 x 8

OpenStudy (cggurumanjunath):

-4=(-15/8)+C -4+(15/8)=C

OpenStudy (cggurumanjunath):

(-32+15)/8=C

OpenStudy (cggurumanjunath):

((-32+15)/8)=C -17/8 =C

OpenStudy (cggurumanjunath):

IS IT CLEAR @Power2Knowledge ?

OpenStudy (cggurumanjunath):

y=mx+c--- standard form the equation of the given line BECOMES-FINAL ANSWER IS BELOW ! Y=(-3/8)X-17/8 Y=(-3X-17)/8

OpenStudy (cggurumanjunath):

Y=(-3/8)X-17/8 FINAL ANSWER